OpenStudy (anonymous):

Graph y=(e^x)(sinx), find zeros, max/mins, and inflections points between or equal to -pi and +pi

OpenStudy (anonymous):

to find the max/min, you basically have to equate the first derivative to zero. that will give you a value of "x". now after having done that, find the second derivative. if it is grater than zero, it will have min. value, if it is less than zero, it will have max. value

OpenStudy (anonymous):

@1ace do you get it?

OpenStudy (anonymous):

ok so i got y' = e^x(sinx+cosx) but i dont understand how to get max/ min points for that one

OpenStudy (anonymous):

now equate that to zero.

OpenStudy (anonymous):

e^x cannot be zero. so, sinx = - cosx tanx = -1 x= tan^-1(-1)

OpenStudy (anonymous):

!! oh wow i never saw that! ok so x= 3pi/4 and 7pi/4 ?

OpenStudy (anonymous):

only 3pi/4. 7pi/4 is out of the given limits

OpenStudy (anonymous):

ooh ok so for the second derivative y''= e^xsinx + e^xcosx - e^xsinx + e^xcosx ? could i cancel out the sinx's there?

OpenStudy (anonymous):

so for this one i would find cosx=0 and x= pi/2 and -pi/2 ?

OpenStudy (anonymous):

so the given function has a min. value

OpenStudy (anonymous):

no. the second derivative is greater than zero, so the function has a min. value

OpenStudy (anonymous):

you don't equate the second derivative to zero.

OpenStudy (anonymous):

but dont i equate y'' to zero to find inflection points?

OpenStudy (anonymous):

fro the point of inflection, yes.

OpenStudy (anonymous):

*for

OpenStudy (anonymous):

right so y''= e^x(2cosx) = 0 where cosx = 0, then x= pi/2 and -pi/2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so in total i have 4 points max/min and inflection points i think i get this question now. Thanks !!

OpenStudy (anonymous):

no problem